# Can LLMs Grasp Implicit Cultural Values? Benchmarking LLMs’ Metacognitive Cultural Intelligence with CQ-Bench

### v1

@@ Abstract
Cultural Intelligence (CQ) refers to the ability to understand unfamiliar cultural contexts—a crucial skill for large language models (LLMs) to effectively engage with globally diverse users. While existing research often focuses on explicitly stated cultural norms, such approaches fail to capture the subtle, implicit values that underlie real-world conversations. To address this gap, we introduce CQ-Bench, a benchmark specifically designed to assess LLMs’ capability to infer implicit cultural values from natural conversational contexts. We generate multi-character conversation-based stories dataset using values from the World Value Survey and GlobalOpinions datasets, with topics including ethical, religious, social, and political. Our dataset construction pipeline includes rigorous validation procedures—incorporation, consistency, and implicitness checks—using GPT-4o, with 98.2% human-model agreement in the final validation. Our benchmark consists of three tasks of increasing complexity: attitude detection, value selection, and value extraction. We find that while o1 and Deepseek-R1 models reach human-level performance in value selection (0.809 and 0.814), they still fall short in nuanced attitude detection, with F1 scores of 0.622 and 0.635, respectively. In the value extraction task, GPT-4o-mini and o3-mini score 0.602 and 0.598, highlighting the difficulty of open-ended cultural reasoning. Notably, fine-tuning smaller models (e.g., LLaMA-3.2-3B) on only 500 culturally rich examples improves performance by over 10% overall, even outperforming stronger baselines (o3-mini) in some cases. Using CQ-Bench [^1] The code and the dataset are available at https://github.com/uscnlp-lime/CQ-Bench, we provide insights into the current challenges in LLMs’ CQ research and suggest practical pathways for enhancing LLMs’ cross-cultural reasoning abilities.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in understanding and generating culturally relevant text (Li et al., 2024a; c; b; Putri et al., 2024). Prior research on cultural alignment in LLMs primarily focuses on modeling differences between national cultures or aligning model outputs with culturally specific norms (Pujari & Goldwasser, 2024; Kharchenko et al., 2024; Shi et al., 2024; Wang et al., 2024; Zhong et al., 2024; Rozen et al., 2024; Johnson et al., 2022). However, in real-world interactions, cultural values are not solely determined by demographic characteristics. Individuals within the same cultural group can hold diverse, nuanced beliefs (Fischer & Poortinga, 2012), and LLMs risk misinterpreting human perspectives if they rely solely on broad demographic generalizations. This limitation becomes particularly evident in human-AI interactions, where effective communication relies not only on an LLM’s ability to recognize explicit cultural markers but also on its capacity to infer implicit cultural values. In human-AI interactions, personalization and customization are increasingly crucial. While LLMs can produce diverse responses when a persona and values are explicitly provided (Grassi et al., 2024), their effectiveness also depends on their ability to infer and align with human values even when such information is not directly given.
Cultural intelligence (CQ) refers to an outsider’s ability to interpret unfamiliar and ambiguous cultural cues naturally (Blanchard & Mohammed, 2024; Earley & Ang, 2003). For LLMs, CQ is crucial for engaging in meaningful conversations with individuals from diverse backgrounds. Unlike traditional assessments of value understanding that focus on detecting explicit statements about cultural norms (Ren et al., 2024; Kiesel et al., 2023), real-life interactions require deeper contextual reasoning—understanding implicit beliefs embedded in everyday speech and actions. Humans do not typically express their values in a structured debate format; instead, values are subtly embedded into casual conversations and personal anecdotes. To enable LLMs to navigate these complexities, we need a more robust approach to evaluating and improving their ability to understand cultural values without relying on explicit cultural knowledge.
@@
In this work, we introduce CQ-Bench, a benchmark designed to evaluate LLMs’ ability to infer implicit cultural values within conversational contexts (Figure 1). We propose a structured methodology for generating multi-turn, multi-character conversation that naturally embeds cultural values. To ensure dataset quality, we perform automatic incorporation, consistency, and implicitness checks using GPT-4o. Human evaluation shows a 98.2% agreement with the model’s validation. We evaluate various LLMs on three increasingly challenging tasks—attitude detection, value selection, and value extraction—to measure their ability to recognize, interpret, and extract cultural values from conversations. While larger models approach human-level performance, smaller models struggle with cultural reasoning without fine-tuning. Performance varies by category: models extract political values well (above 0.7 $F_{1}$ score overall) but worse in religious values (below 0.60). Notably, we find that fine-tuning on just 500 data points significantly enhances smaller models’ ability to detect out-of-domain cultural values, demonstrating the efficiency to improve LLMs’ capabilities with minimal data. Our contributions in this work can be summarized as:
@@
  • Benchmark Design: We introduce CQ-Bench, a novel benchmark for evaluating LLMs’ cultural intelligence by assessing their ability to infer implicit cultural values from conversational contexts.
  • Comprehensive Evaluation: We construct three escalating tasks—attitude detection, value selection, and value extraction—to systematically evaluate LLMs’ cultural reasoning across multiple domains and model scales.
  • Data-Efficient Enhancement: We demonstrate that fine-tuning with just 500 culturally rich examples significantly improves performance on smaller models, highlighting an efficient strategy for boosting cultural intelligence.
2 CQ-Bench Design and Implementation
@@
2.1 Data Generation
Our goal is to generate conversation-based stories featuring 4–5 characters, with cultural values implicitly embedded in the narrative. We begin by outlining the process of value selection, followed by the story generation approach. Finally, we describe the validation steps taken to ensure the quality of the resulting dataset.
Cultural Value Selection
Cultural values are defined as values inherently linked to culture and expressed through distinct attitudes. Each cultural value consists of two components: (1) Statement, which presents or solicits an opinion, and (2) Attitude, which signifies agreement or disagreement with the statement. As shown in Figure 1, “The only acceptable religion is my religion” is a statement while “Disagree” is an attitude. Different statements offer multiple attitude options, aligning with the settings in the original questionnaires.
The statements in this study are sourced from the World Values Survey (WVS) (Haerpfer et al., 2022) and the GlobalOpinion dataset (Durmus et al., 2023). The WVS is a global research project that explores individuals’ values and beliefs, how their evolution over time, and their sociopolitical implications. We manually select values that either focus on personal beliefs or characterize societal and community attributes. The GlobalOpinion dataset contains a subset of survey questions about global issues and public opinions, adapted from the WVS and the Pew Global Attitudes Survey.
In CQ-Bench, each story incorporates five cultural values randomly selected from a list generated in three distinct settings:
  • Random Setting: Aligning with previous research (Li et al., 2024a), which utilizes a subset of 50 values from the WVS, we randomly select 5 statements from this subset and assign each a random attitude to form cultural values.
  • Multiple Attitude Setting: The first two settings require consistency in attitude reflection throughout the story. In contrast, this setting assigns each character a distinct attitude toward the selected values, fostering diverse perspectives and enhancing complexity. For example, in the first two settings, one value could be Work is a duty towards the society – agree. In the multiple attitude setting, the values could be Alice: Work is a duty towards the society – agree and Raj: Work is a duty towards the society – disagree.
Story Generation
Our goal is to create a dataset of conversation-based stories that reflect cultural values. Different from Li et al. (2024b), where characters sequentially present their ideas in a debate format, conversations in our CQ-Bench adopt a more natural and casual real-life style. Each story is generated based on five specified cultural values and a predefined scenario setting (Appendix A.2). The generated stories must adhere to the following guidelines:
@@ 2.1 Data Generation
Post-Generation Checking
All generated stories undergo a post-check to ensure quality. The checklist corresponds to the first three guidelines, while the fourth is verified directly using word count. Through preliminary experiments, we find that GPT-4o is better in verification while GPT-4o-mini can perform as well as GPT-4o in refining.
@@
  • Consistency Check: In both the Random and Category-Specific settings, all characters in the story are expected to adhere to the assigned value. However, in some cases, certain characters may express opposing views, particularly when the assigned value contradicts prevailing social norms. For instance, if the given value is “If women go to work, the children will suffer—agree,” a character might challenge this stance. GPT-4o is used to check consistency. When inconsistencies are detected, we revise the conflicting speech to align with the assigned value. Specifically, we provide GPT-4o with the full story, along with the original speech, and prompt it to generate only the revised version that conforms to the intended value. We do not use GPT-4o-mini for revision, as we found it is unable to effectively rewrite inconsistent speech and tends to follow the original stance.
@@
To enhance reliability, we perform three rounds of incorporation and consistency checks as final validation. Missing and contradictory values are addressed using a majority vote approach based on three evaluations. Specifically, missing values are removed from the ground truth, while contradictory values are documented along with the specific inconsistencies and the characters exhibiting inconsistent speech. Documented contradictory values are used in generating dataset. More details are shown in Appendix A.2.
In particular, resolving inconsistencies in speech rewriting remains challenging, even when the model successfully detects contradictions. However, we find that detecting these issues is generally easier for the model than correcting them. We conduct human validation for incorporation and consistency checks on 50 samples, comparing the results with model validation. Using Cohen’s kappa, we calculate an agreement score of 0.658 for the incorporation check. To further analyze alignment, we remove instances where the model assigned a label of 0 (not incorporated), ensuring that corresponding values are also excluded from ground truth values. Among the remaining cases, where the model consistently identifies incorporation, 98.2% align with human judgment, demonstrating the faithfulness of the model’s validation. Full agreement scores are provided in Appendix A.3.
2.2 Task Definition
As shown in Figure 1, CQ-Bench includes three tasks with increasing difficulties: attitude detection, value selection, and value extraction.
Attitude Detection (AD)
Given a story $S$ and a statement $T$, the model identifies the attitude expressed within $S$. In the Random and Category-Specific settings, it determines the overall attitude including all characters, excluding cases where contradictions arise. If a story contains inconsistent mentions of certain values, those values will be removed when constructing the dataset. In the Multiple Attitude Setting, the model identifies the attitude of a specific character. The model selects an answer from a predefined set of options, making this a multiple-choice task.
Value Selection (VS)
Given a story $S$ and a predefined set of 15 candidate values $\mathcal{V}=\{v_{1},v_{2},\dots,v_{15}\}$, the model is required to select exactly $X$ ground-truth values, denoted as $\mathcal{V}^{*}\subset\mathcal{V}$, where $|\mathcal{V}^{*}|=X$. The remaining $15-X$ options are randomly sampled from non-ground truth values. This task is more challenging than attitude detection, as the model must first identify the relevant topics before selecting the correct values. Formally, the model must learn a function $f(S,\mathcal{V})\rightarrow\mathcal{V}^{*}$, where $\mathcal{V}^{*}\subset\mathcal{V}$ and $|\mathcal{V}^{*}|=X$.
Value Extraction (VE)
Given a story $S$, the model is asked to extract key cultural values on given topics without predefined answer choices (e.g., social, ethical, political, etc.). The model is provided with examples illustrating the expected format of values, along with a set of topics $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ that cover all seed values in the prompt. For each topic $t_{i}\in\mathcal{T}$, the model generates a set of values $\mathcal{V}_{t_{i}}$, where:
$\mathcal{V}_{t_{i}}=\begin{cases}\{v_{1},v_{2},\dots,v_{m}\},&\text{if relevant values exist}\\[8.0pt] \emptyset,&\text{otherwise}\end{cases}.$
While the model is encouraged to provide comprehensive and detailed responses during the reasoning phase, we ask it to limit the answer size to 10 total values for easier performance comparison across models. Evaluation is based on recall, measuring the proportion of ground-truth values correctly identified.
@@ 3 Evaluating LLMs with CQ-Bench
In this section, we focus on three research questions (RQs):
  • RQ1.
    How does cultural intelligence vary across different LLMs?
  • RQ2.
    Does LLMs’ cultural intelligence show discrepancy over different topics?
  • RQ3.
    How can we improve models’ cultural intelligence?
@@ 3.1 Experimental Settings
Models Selection
We select a diverse range of models, including both open-source and closed-source models, with varying sizes. For open-source models, we include Qwen 2.5 (7B, 14B, and 32B) (Bai et al., 2023), LLaMA 3.1 8B, and LLaMA 3.2 3B (Grattafiori et al., 2024), Deepseek-V3 and Deepseek-R1 (DeepSeek-AI, 2025). Additionally, we experiment with DeepSeek-Distill models (Qwen 2.5 1.5B, Qwen 2.5 7B, LLaMA 3.1 8B), which have been reported to achieve remarkable performance on math and coding tasks, even at smaller scales. Our goal is to evaluate if they can also perform well on culture-related reasoning tasks. Finally, for closed-source models, we select GPT-4o-mini, o3-mini, and o1 (Jaech et al., 2024). We also conduct analyses with three human participants who complete the same tasks on 25 stories, which took 5-6 hours in total.
Evaluation Metrics
For attitude detection and value selection, we use the F1 score to compare the predicted answers with the ground truth values. For value extraction, which is more open, we employ LLM-as a judge (Zheng et al., 2023) to evaluate the responses. Let $V=\{v_{1},v_{2},\dots,v_{n}\}$ be the set of ground truth values, and let $\hat{V}=\{\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{m}\}$ be the set of predicted values. We use GPT-4o to access the output. We ask humans to conduct the same evaluation as GPT-4o on 25 stories. The agreement score is 0.864, which shows the reliability of LLM-as-a-judge.
For each ground truth value $v_{i}$, we define the score function as follows:
$S(v_{i},\hat{V})=\begin{cases}1,&\text{if }v_{i}\text{ is fully presented in }\hat{V}\\ 0.5,&\text{if }v_{i}\text{ is partially presented in }\hat{V}\\ 0,&\text{if }v_{i}\text{ is not mentioned in }\hat{V}\end{cases}$
@@
Dataset
We generate 500 stories for the random setting, 100 stories for each category-specific setting, and 100 stories for the multiple-attitude setting. For story generation, we use GPT-4o-mini, while GPT-4o is used for validation. In the multiple-attitude setting, we implement only the attitude detection task. The value selection/extraction task is not included because the ground truth values typically exceed ten, making it challenging for the model to choose accurately. We report the total datapoints for each task in Table 4.
Summarize-then-analyse long CoT prompting
For attitude detection and value selection, we experiment with two settings: no reasoning and CoT reasoning. In the no reasoning setting, we ask models to provide answers directly without explanation. In the reasoning setting, we provide a step-by-step reasoning guideline to guide the models’ responses. For attitude detection (AD), we instruct models to first summarize speech relevant to the given statement and then analyze the attitude based on the retrieved speech.
For value selection (VS), we employ a multi-step approach: (1) The model summarizes the topics mentioned in the story based on the provided options; (2) Selects values associated with the identified topics. (3) Reasons about which value best reflects the story. (4) Finally, outputs the selected value. For value extraction (VE), we ask the model to summarize the content for each topic and then predict relevant values based on the summarization.
3.2 Cultural Intelligence across Different LLMs
We first show the results of attitude detection and value selection in Table 1. Overall, larger models outperform smaller models by a lot and Summarize-then-analyse prompting can improve performance generally. Human participants achieve an average score of 0.689 and 0.768 on AD and VS respectively.
@@ 3.2 Cultural Intelligence across Different LLMs
7B 14B 32B 8B 3B Q 1.5B Q 7B L 8B 4o-mini o3-mini o1 V3 R1
AD W/O R 0.529 0.553 0.572 0.527 0.455 - - - 0.604 0.622 0.622 0.642 0.595
W/ R 0.620 0.616 0.624 0.506 0.372 0.381 0.484 0.556 0.639 0.661 0.622 0.660 0.635
Merged 0.783 0.778 0.786 0.631 0.480 0.490 0.622 0.705 0.820 0.793 0.811 0.837 0.837
Multiple 0.592 0.621 0.642 0.590 0.462 0.403 0.510 0.584 0.645 0.684 - 0.691 -
VS W/O R 0.515 0.585 0.633 0.421 0.272 - - - 0.639 0.759 0.810 0.780 0.798
W/ R 0.374 0.607 0.717 0.274 0.1 0.383 0.411 0.418 0.576 0.779 0.809 0.819 0.814
VE - - 0.629 - - - - - 0.602 0.598 0.621 0.704 0.736
Table 1: Results on Attitude Detection (AD) and Value Selection (VS). For deepseek-distill models, the models always output their thinking process i.e. reasoning. Therefore, we only report reasoning results for those models “W/ R“. For the o1 model, we only report results on the random setting dataset. In the “Merged” setting, we merge similar options and in “Multiple” setting, story characters may hold different opinions.

### v2

# Can LLMs Grasp Implicit Cultural Values? Benchmarking LLMs’ Cultural Intelligence with CQ-Bench

@@ Abstract
Cultural Intelligence (CQ) refers to the ability to understand unfamiliar cultural contexts—a crucial skill for large language models (LLMs) to effectively engage with globally diverse users. Existing studies often focus on explicitly stated cultural norms, but fail to capture the subtle, implicit values that are common in daily conversation. To address this gap, we introduce CQ-Bench, a benchmark specifically designed to assess LLMs’ capability to infer implicit cultural values from natural conversational contexts. CQ-Bench consists of multi-character conversation-based stories using values from the World Value Survey and the GlobalOpinions, with topics including ethical, religious, social, etc. Our automatic dataset construction pipeline integrates rigorous validation procedures (incorporation, consistency, and implicitness checks), achieving a 94.5% human–model agreement in the final validation. To leverage CQ-Bench data, we design three tasks of increasing complexity: attitude detection, value selection, and value extraction. These tasks evaluate whether models can detect attitude and recognize values embedded within natural dialogues rather than relying on explicit cultural knowledge. We find that while frontier models like o1 reach human-level performance in value selection (0.809 $F_{1}$), they still fall short in nuanced attitude detection (0.622 $F_{1}$). Notably, fine-tuning a smaller LLaMA-3.2-3B on only 500 culturally-rich examples improves performance by over 10%, even outperforming o3-mini in some cases. Using CQ-Bench, we provide insights into the current challenges in LLMs’ CQ research and suggest practical pathways for enhancing LLMs’ cross-cultural reasoning abilities.  [^1] The code is available at: https://github.com/uscnlp-lime/CQ-Bench
Can LLMs Grasp Implicit Cultural Values? Benchmarking LLMs’ Cultural Intelligence with CQ-Bench
Ziyi Liu1    Priyanka Dey1    Jen-tse Huang3      Zhenyu Zhao1    Bowen Jiang4 Rahul Gupta2    Yang Liu2   Yao Du1   Jieyu Zhao1 1University of Southern California    2Amazon AGI 3John Hopkins University  4University of Pennsylvania {zliu2803, deyp, zzhao767, jieyuz}@usc.edu {gupra, yangliud}@amazon.com jhuan236@jh.edu bwjiang@seas.upenn.edu
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in understanding and generating culturally relevant text (Li et al., 2024a, c, b; Putri et al., 2024). Prior research on LLM cultural alignment primarily focuses on modeling differences between national cultures (Pujari and Goldwasser, 2024; Kharchenko et al., 2024; Shi et al., 2024; Wang et al., 2024) or aligning models with culturally specific norms (Zhong et al., 2024; Rozen et al., 2024; Johnson et al., 2022; Kim et al., 2024; Monazzah et al., 2025).
@@
However, cultural values in real-world interactions are not solely determined by demographic characteristics. Individuals within the same cultural group can hold diverse and even conflicting beliefs (Fischer and Poortinga, 2012), and LLMs risk oversimplifying human perspectives if they rely merely on broad demographic generalizations Saha et al. (2025a). This limitation becomes particularly problematic in human-AI interactions, where successful communication relies not only on an LLM’s ability to recognize explicit cultural markers but also on its capacity to infer implicit cultural values. While LLMs can produce diverse responses when a persona and values are explicitly provided (Grassi et al., 2024; Wang et al., 2025; Huang et al., 2024; Jiang et al., 2025), real-world communication often lacks such explicit signals. Humans do not typically express their values in a structured debate format Chatterji et al. (2025); instead, values are subtly embedded into casual conversations and personal anecdotes.
Cultural intelligence (CQ) refers to an outsider’s ability to interpret unfamiliar and ambiguous cultural cues (Blanchard and Mohammed, 2024; Earley and Ang, 2003). For LLMs, CQ is crucial for engaging in meaningful conversations with individuals from diverse backgrounds. Unlike traditional assessments of value understanding that focus on detecting explicit statements about cultural norms (Ren et al., 2024; Kiesel et al., 2023), it emphasizes deeper contextual reasoning—mirroring real-life interactions where implicit beliefs are embedded in everyday speech and action.
@@
To address these gaps, we introduce CQ-Bench, a benchmark designed to evaluate LLMs’ ability to infer implicit cultural values within conversational contexts (Figure 1). We propose a structured automatic pipeline for generating multi-turn, multi-character conversations that naturally embed cultural values. We design additional incorporation, consistency, and implicitness checks using GPT-4o as a judge to ensure dataset quality. Human evaluation shows a 94.5% agreement with the GPT-4o’s judgment. We then evaluate various LLMs on three increasingly challenging tasks based on CQ-Bench: attitude detection, value selection, and value extraction, to measure their ability to recognize, interpret, and extract cultural values from conversations. While larger models generally show more promising results than smaller models in most tasks, our detailed analysis reveals notable performance variations across different categories of values. For example, models extract political values well (above 0.7 $F_{1}$ score overall) but perform much worse in religious values (below 0.6), highlighting substantial room for improving LLMs’ understanding of culture values.
Our contributions are as follows:
@@
  • Benchmark Design: We are the first to benchmark LLMs’ cultural intelligence via CQ-Bench, which assesses their ability to infer implicit cultural values in conversation.
  • Automatic Dataset Generation Pipeline: We introduce a pipeline that automatically generates high-quality synthetic datasets. The strong agreement between our model and human evaluations demonstrates the robustness of the pipeline and ensures reproducibility for future research.
  • Comprehensive Evaluation: We construct three escalating tasks—attitude detection, value selection, and value extraction—to systematically evaluate open-sourced and close-sourced models’ CQ. We further perform both supervised and reinforcement fine-tuning with GRPO to enhance the reasoning and alignment.
2 CQ-Bench
Evaluating a model’s cultural intelligence goes beyond simply identifying a broad value in a short speech Ren et al. (2024). To reflect real-world scenarios, we design comprehensive tasks that examine implicit cultural value understanding from multiple angles. As shown in Figure 1, CQ-Bench features three tasks of growing difficulty: attitude detection, value selection, and value extraction. We will first introduce the definition of culture value and conversation setup in CQ-Bench and then explain the tasks in detail.
2.1 Culture Value
Cultural values are defined as values inherently linked to culture and expressed through different attitudes. Each cultural value consists of two components: (1) Statement, which presents or solicits an opinion, and (2) Attitude, which signifies agreement or disagreement with the statement. As shown in Figure 1, “The only acceptable religion is my religion” is a statement while “Disagree” is an attitude. Different statements offer multiple attitude options, aligning with the settings in the original questionnaires. The statements in this study are sourced from the World Values Survey (WVS) (Haerpfer et al., 2022) and the GlobalOpinion dataset (Durmus et al., 2023). The WVS is a global research project that explores individuals’ values and beliefs, how their evolution over time, and their sociopolitical implications. The GlobalOpinion dataset contains a subset of survey questions about global issues and public opinions, adapted from the WVS and the Pew Global Attitudes Survey. We manually select values that either focus on personal beliefs or characterize societal and community attributes.
2.2 Conversation Setting
In CQ-Bench, each conversation story features 4–5 characters, with randomly selected cultural values implicitly embedded in the narrative. We consider three distinct settings:
  • Random Setting: Aligning with previous research (Li et al., 2024a), which utilizes a subset of 50 values from the WVS, we randomly select 5 statements from this subset and assign each a random attitude shared by all characters to form cultural values.
  • Multiple Attitude Setting: In contrast to first two settings, this setting assigns each character a distinct attitude toward the selected values, fostering diverse perspectives and enhancing complexity. For example, in the first two settings, one value could be Work is a duty towards the society – agree. In the multiple attitude setting, the values could be Alice: Work is a duty towards the society – agree and Raj: Work is a duty towards the society – disagree.
2.3 Attitude Detection (AD)
To understand people’s values, the first step is to examine whether models can recognize preferences on specific topics. Accordingly, we evaluate whether the model can interpret attitudes expressed in conversation. Given a story $S$, a statement $T$, and a limited set of attitude options $\mathcal{O}=\{o_{1},o_{2},\dots,o_{n}\}$ (e.g., Agree, Disagree, Very often, Not at all often), the model identifies the attitude toward $T$ expressed in $S$ from options $\mathcal{O}$. In the Random and Category-Specific settings, it determines the overall attitude including all characters. In the Multiple Attitude setting, characters may hold different stances, and the model must identify the attitude of a specified character.
2.4 Value Selection (VS)
Given a story $S$ and a predefined set of 15 candidate values $\mathcal{V}=\{v_{1},v_{2},\dots,v_{15}\}$ (candidate values are different for each datapoint), the model is required to select exactly $X$ ground-truth values, denoted as $\mathcal{V}^{*}\subset\mathcal{V}$, where $|\mathcal{V}^{*}|=X$. The remaining 15X15-X options are randomly sampled from non-ground truth values. This task is more challenging than attitude detection, as the model must first identify the relevant topics before selecting the correct values. Formally, the model must learn a function $f(S,\mathcal{V})\rightarrow\mathcal{V}^{*}$, where $\mathcal{V}^{*}\subset\mathcal{V}$ and $|\mathcal{V}^{*}|=X$.
2.5 Value Extraction (VE)
In real-world scenarios, there is no predefined set of values for models to select from, making value understanding particularly challenging. To address this, we design Value Extraction to assess cultural value detection without prior knowledge. Given a story $S$, the model extracts key cultural values across given topics (e.g., social, ethical, political) without predefined choices. It receives examples of expected formats and a topic set $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ covering all seed values. For each topic $t_{i}\in\mathcal{T}$, the model outputs a value set $\mathcal{V}_{t_{i}}=\{v_{1},v_{2},\dots,v_{m}\}$ if relevant values exist, or $\mathcal{V}_{t_{i}}=\emptyset$ otherwise.
We ask the model to limit the answer size to 10 total values for easier performance comparison across models. Evaluation is based on recall, measuring the proportion of ground-truth values correctly identified.
3 Automatic Dataset Generation
@@
Our goal is to build a dataset of conversation-based stories that reflect cultural values. Unlike Li et al. (2024b), where characters present their ideas sequentially in a debate format, conversations in CQ-Bench are designed to be more natural and casual, resembling real-life interactions. However, no existing dataset captures real human conversations in cultural contexts, and collecting such data is difficult due to the extensive human labeling required. Following prior work Ying et al. (2025); Sui et al. (2025), we therefore construct a synthetic dataset using an automatic pipeline that both generates and validates the data.
3.1 Generation guidelines
Each story is generated based on five specified cultural values and different scenarios for diversity (Appendix A.2). The generated stories must adhere to the following guidelines:
@@ 3.1 Generation guidelines
3.2 Post-Generation Checking
All generated stories undergo a post-check to ensure quality, as shown in Figure 2. The checklist corresponds to the first three guidelines, while the fourth is verified directly using word count. Through preliminary experiments, we find that GPT-4o is better in verification while GPT-4o-mini can perform as well as GPT-4o in refining.
@@ 3.2 Post-Generation Checking
  • Consistency Check: In both the Random and Category-Specific settings, all characters in the story are expected to adhere to the assigned value. However, in some cases, certain characters may express opposing views, particularly when the assigned value contradicts prevailing social norms. When inconsistencies are detected, we revise the conflicting speech to align with the assigned value. Specifically, we provide GPT-4o with the full story, along with the original speech, and prompt it to generate only the revised version that conforms to the intended value. We do not use GPT-4o-mini for revision, as we found it is unable to effectively rewrite inconsistent speech and tends to follow the original stance.
@@
To enhance reliability, we perform three rounds of incorporation and consistency checks as final validation. Missing and contradictory values are addressed using a majority vote approach based on three evaluations. Specifically, missing values are removed from the ground truth, while contradictory values are documented along with the specific inconsistencies and the characters exhibiting inconsistent speech. Documented contradictory values are used in generating dataset (Appendix A.2).
Human evaluation
We conduct human validation to assess each step of the conversation refinement process and the overall conversation quality. A total of 150 conversations, encompassing 750 values, are evaluated across four dimensions: value reflection, consistency, implicitness, and naturalness. For value reflection and consistency, we compare human annotations with model predictions. For implicitness, annotators compare each speech segment before and after refinement and rate the degree of improvement on a 1–5 scale. For naturalness, they rate how natural the refined conversation sounds, also on a 1–5 scale. Fifty computer science students participated as volunteers, each labeling a small subset of the data, with every data point annotated by two independent annotators. The Cohen’s $\kappa$ score for value reflection is 0.546, indicating moderate agreement. After removing instances where the model judged a value as not reflected (and the corresponding ground-truth values), the remaining cases show 94.5% alignment with human judgment, demonstrating the faithfulness of the model’s validation. Similarly, the consistency validation achieves a Cohen’s $\kappa$ of 0.515 and 92.5% alignment with human annotations. For implicitness refinement, over 80% of cases are rated as significantly improved, with an average improvement score of 3.28. For naturalness, the average rating is 3.63, indicating that most conversations are perceived as natural.
By introducing this automatic pipeline, future work can build upon to efficiently generate high-quality datasets. In particular, the pipeline makes it possible to incorporate a wide range of cultural values, enabling scalable extensions and more diverse applications in value understanding research.
@@ 4.1 Experimental Settings
Models selection.
We select a diverse range of models, including both open-source and closed-source models, with varying sizes. For open-source models, we include Qwen 2.5 (7B, 14B, and 32B), Qwen 3 4B (Bai et al., 2023), LLaMA 3.1 8B, and LLaMA 3.2 3B (Grattafiori et al., 2024), Deepseek-V3 and Deepseek-R1 (DeepSeek-AI, 2025). Additionally, we experiment with DeepSeek-Distill models (Qwen 2.5 1.5B, Qwen 2.5 7B, LLaMA 3.1 8B), which have been reported to achieve remarkable performance on math and coding tasks, even at smaller scales. Our goal is to evaluate if they can also perform well on culture-related reasoning tasks. Finally, for closed-source models, we select GPT-4o-mini, o4-mini, o3, o3-mini, and o1 (Jaech et al., 2024). We also conduct analyses with four human participants who complete the same tasks on 25 stories, which took 5-6 hours in total  [^2] Hourly payment for annotators in this work is \$16.5. .
Evaluation metrics.
For attitude detection and value selection, we use the $F_{1}$ score to compare the predicted answers with the ground truth values. For value extraction, which is more open, we employ LLM-as-a-judge (Zheng et al., 2023) to evaluate the responses. Let $V=\{v_{1},v_{2},\dots,v_{n}\}$ be the set of ground truth values, and let $\hat{V}=\{\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{m}\}$ be the set of predicted values. We use GPT-4o to access the output. We ask humans to conduct the same evaluation as GPT-4o on 25 stories. The agreement score is 0.864, which shows the reliability of LLM-as-a-judge. For each ground truth value $v_{i}$, we define the score function as follows:
[…]
@@
Dataset.
We generate 500 stories for the random setting, 100 stories for each category-specific setting, and 100 stories for the multiple-attitude setting, resulting in a total of 1,000 stories. For story generation, we use GPT-4o-mini, while GPT-4o is used for validation. In the multiple-attitude setting, we implement only the attitude detection task. The value selection/extraction task is not included because the ground truth values typically exceed ten, making it challenging for the model to choose accurately. We report the total datapoints of attitude detection and values covered in the story of value selection for each task in Table 1.
Random Political Social Religious Ethical Multiple
AD 1665 301 213 425 270 1540
VS 2099 402 285 335 351 -
Table 1: Total datapoints for attitude detection (AD) and total values for value selection (VS) tasks.
@@
4B 7B 14B 32B 8B 3B Q 1.5B Q 7B L 8B 4o-mini o3-mini o1 o4-mini o3 V3 R1
AD W/O R 0.560 0.529 0.553 0.572 0.527 0.455 - - - 0.604 0.622 0.622 0.60 0.689 0.642 0.595
W/ R 0.610 0.620 0.616 0.624 0.506 0.372 0.381 0.484 0.556 0.639 0.661 0.622 0.595 0.622 0.660 0.635
Merged 0.775 0.783 0.778 0.786 0.631 0.480 0.490 0.622 0.705 0.820 0.793 0.811 0.834 0.824 0.837 0.837
Multiple 0.607 0.592 0.621 0.642 0.590 0.462 0.403 0.510 0.584 0.645 0.684 - 0.738 - 0.691 -
VS W/O R 0.653 0.515 0.585 0.633 0.421 0.272 - - - 0.639 0.759 0.810 0.828 0.830 0.780 0.798
W/ R 0.468 0.374 0.607 0.717 0.274 0.1 0.383 0.411 0.418 0.576 0.779 0.809 0.820 0.710 0.819 0.814
VE - - - 0.629 - - - - - 0.602 0.598 0.610 0.696 0.732 0.704 0.736
Table 2: Results on Attitude Detection (AD) and Value Selection (VS). For deepseek-distill models, the models always output their thinking process (i.e. reasoning). Therefore, we only report reasoning results for those models “W/ R”. For larger models like Deepseek-R1, o1, and o3, we report results on the same subset used for human evaluation, due to the high cost of running them on the full dataset. In the “Merged” setting, we merge similar options and in “Multiple” setting, story characters may hold different opinions. The “Merged” and the “Multiple” settings are both under reasoning settings.
4.2 Prompt Setup
We implement two prompting strategies: with reasoning and without reasoning. In the no-reasoning setting, models provide answers directly without explanations. In the reasoning setting, we experiment with both zero-shot and few-shot prompting. Preliminary results show that few-shot prompting does not improve performance and, sometimes biases the model toward the values presented in the demonstrations. Therefore we restrict our main experiments to zero-shot reasoning across all models.
Summarize-then-analyse long CoT prompting.
In the reasoning setting, we observe that simple prompts often lead to low scores, sometimes even lower than the no-reasoning baseline. To address this, we provide a step-by-step reasoning guideline to guide the models’ responses. For attitude detection (AD), we instruct models to first summarize speech relevant to the given statement and then analyze the attitude based on the retrieved speech. For value selection (VS), we employ a multi-step approach: (1) The model summarizes the topics mentioned in the story based on the provided options; (2) Selects values associated with the identified topics; (3) Reasons about which value best reflects the story; (4) Finally, outputs the selected value. For value extraction (VE), we ask the model to summarize the content for each topic and then predict relevant values based on the summarization.
4.3 CQ across Different LLMs
We first show the results of attitude detection and value selection in Table 2. Overall, larger models outperform smaller models by a lot and summarize-then-analyse prompting can improve performance generally. Human participants achieve an average score of 0.689 and 0.765 on AD and VS respectively.